OpenStudy (amenah8):

find the extreme values AND where they occur: y = sqrt(-x^2 + 2x +3)

OpenStudy (solomonzelman):

You need to find the critical numbers

OpenStudy (solomonzelman):

y'=?

OpenStudy (solomonzelman):

3 cases when critical numbers occur (1) The x-values that are a solution to 0=f'(x) are critical numbers. (2) The x-values at which f'(x) is undefined, provided f(x) is defined for them. (3) Any closed boundaries are automatically critical numbers.

OpenStudy (solomonzelman):

If want, algebra \(\color{#000000 }{ \displaystyle y=\sqrt{-x^2+2x+3} }\) \(\color{#000000 }{ \displaystyle y=\sqrt{-x^2+2x-1+4} }\) \(\color{#000000 }{ \displaystyle y=\sqrt{-(x^2-2x+1)+4} }\) \(\color{#000000 }{ \displaystyle y=\sqrt{-(x-1)^2+4} }\) \(\color{#000000 }{ \displaystyle y^2=-(x-1)^2+4}\) circle centered at (1,0) with radius of 4

OpenStudy (solomonzelman):

well, this is really the upper half of the circle, because your range was \(y\ge0\)

OpenStudy (solomonzelman):

and my bad, the radius squared is 4, but radius just is 2.

OpenStudy (solomonzelman):

so the minimums are (1±2,0) ad maximum is up top so the minimums are (1,0+2)

OpenStudy (amenah8):

thank you so much! sorry i was unable to respond, my computer was acting up. but i understand the problem now. thank you!

OpenStudy (solomonzelman):

was it algebra or calculus? (curious)

OpenStudy (amenah8):

calculus. i solved for the derivative and got this: .5(-x^2 + 2x + 3)^-.5 (-2x+2)

OpenStudy (amenah8):

is that correct?

OpenStudy (solomonzelman):

Yes, you are correct! (Good chain rule :) )

OpenStudy (solomonzelman):

And with that you need critical numbers

OpenStudy (solomonzelman):

3 cases when critical numbers occur (1) The x-values that are a solution to 0=f'(x) are critical numbers. (2) The x-values at which f'(x) is undefined, provided f(x) is defined for them. (3) Any closed boundaries are automatically critical numbers.

OpenStudy (amenah8):

from that, without using the algebra you did above, how would i find the min and max?

OpenStudy (solomonzelman):

you aren't given intervals (let's say we didn't know it is a circle)

OpenStudy (amenah8):

i know i have to solve for 0, but i got stuck :(

OpenStudy (solomonzelman):

Yes set f'(x)=0

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 0=\frac{-2x-2}{\sqrt{-x^2+2x+3}} }\)

OpenStudy (solomonzelman):

but wait, it is not corrct

OpenStudy (solomonzelman):

you have a 2 in denominator when u differentiate the root

OpenStudy (solomonzelman):

Am I right, or not?

OpenStudy (amenah8):

why is a 2 in the denominator when you take the derivative?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{d }{dx} x^{1/2}=(1/2)x^{-1/2} }\)

OpenStudy (solomonzelman):

via the power rule... right

OpenStudy (amenah8):

so now i set (1/2)x^-(1/2) equal to 0?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f'(x)=\frac{-2x-2}{2\sqrt{-x^2+2x+3}} }\)

OpenStudy (solomonzelman):

ion our case, correct?

OpenStudy (solomonzelman):

oh, sorry +2

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f'(x)=\frac{-2x+2}{2\sqrt{-x^2+2x+3}} }\)

OpenStudy (amenah8):

and set that equal to 0 and solve for x?

OpenStudy (solomonzelman):

yup

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 0=\frac{-2x+2}{2\sqrt{-x^2+2x+3}} }\)

OpenStudy (solomonzelman):

x=1

OpenStudy (solomonzelman):

and case (2)

OpenStudy (solomonzelman):

where is f(x) defined, but f'(x) NOT defined?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(x)=\sqrt{-x^2+2x+3} }\) \(\color{#000000 }{ \displaystyle f'(x)=\frac{-2x+2}{2\sqrt{-x^2+2x+3}} }\)

OpenStudy (solomonzelman):

(Hint- in derivative unlike in the f(x), the root cannot be =0)

OpenStudy (amenah8):

ah, sorry ignore that^

OpenStudy (amenah8):

no, i meant ignore what i said about setting it equal to 0 :) i accidentally copied a post i'd made earlier :)

OpenStudy (solomonzelman):

:)

OpenStudy (amenah8):

well, the equation wouldn't work if x = 0, because you cannot divide by 0 ... is that what you meant?

OpenStudy (solomonzelman):

no, if the denominator is 0! (but you are thinking correctly)

OpenStudy (solomonzelman):

if x=0, denom is not 0

OpenStudy (solomonzelman):

there are two solutions for denominator=0

OpenStudy (solomonzelman):

(by exclamation mark neaxt to 0, "0!" I don't mean zero factorial :))

OpenStudy (amenah8):

so the denominator = 0 when x = .... well, 2 time the sqrt(0) = 0 how can i set what is in the sqrt to equal 0?

OpenStudy (solomonzelman):

You can simply set what is inside the square root =0, so I don't really understand your question...

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f'(x)=\frac{-2x+2}{2\sqrt{-x^2+2x+3}} }\) -x²+2x+3 =0

OpenStudy (amenah8):

ah! i see.

OpenStudy (amenah8):

and solve for x?

OpenStudy (solomonzelman):

yes.

OpenStudy (amenah8):

one moment while i figure it out please!

OpenStudy (solomonzelman):

sure, take your time../.

OpenStudy (amenah8):

-x^2 + 2x = -3? and then could i combine the -x^2 + 2x to make -2x^3?

OpenStudy (solomonzelman):

WOW

OpenStudy (amenah8):

or are they different roots and so cannot be combined?

OpenStudy (amenah8):

(sorry, algebra is the biggest math concept i struggle with :)

OpenStudy (solomonzelman):

That is very bad!

OpenStudy (solomonzelman):

Algebra is key to "A" in calculus or any other course in math...

OpenStudy (solomonzelman):

they are not "like terms"

OpenStudy (amenah8):

hmmm. that must be why I'm stuck with a B+.... :)

OpenStudy (solomonzelman):

You seem to be good at calculus, but you really need to nail the algebra....

OpenStudy (amenah8):

yeah :)

OpenStudy (solomonzelman):

You need to know the entire algebra (look common course outline)

OpenStudy (solomonzelman):

the solutions to -x²+2x+3 are x=-1 and x=3

OpenStudy (solomonzelman):

but, later, please look and make sure you know absolutely all of it.

OpenStudy (amenah8):

once you had -x^2 + 2x = -3, what did you do to the left side of the equation to get just x=?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle -x^2+2x+3=0}\) \(\color{#000000 }{ \displaystyle x^2-2x=3}\) \(\color{#000000 }{ \displaystyle x^2-2x+1=4}\) \(\color{#000000 }{ \displaystyle (x-1)^2=4}\) \(\color{#000000 }{ \displaystyle x-1=\pm 2}\) \(\color{#000000 }{ \displaystyle x=\pm 2+ 1}\)

OpenStudy (solomonzelman):

completing the square -ma fav. but precisely,

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle (x-1)^2=4}\) when you take the square root of both sides, you get \(\color{#000000 }{ \displaystyle|x-1|=2}\) and that is where we get \(\color{#000000 }{ \displaystyle x-1=\pm 2}\)

OpenStudy (solomonzelman):

this is via the defition of abs value \(\color{#000000 }{ \displaystyle \sqrt{r^2}=|r|}\)

OpenStudy (amenah8):

so are the solutions still x=-1 and x=3? what does the x=±2+1 represent?

OpenStudy (solomonzelman):

\(\pm2\) says that 2 can be neg or pos

OpenStudy (amenah8):

how did you get x = -1 and x=3 from that?

OpenStudy (solomonzelman):

so \(\pm 2+1\) read as, either \(+2+1\) OR \(-2+1\)

OpenStudy (solomonzelman):

3 and -1, respectively

OpenStudy (solomonzelman):

I wonder how you pass the class at all without algebra. I a saying this because I want you to know algebra.

OpenStudy (amenah8):

yeah... :) recently, most of our tests have been finding the derivative, so I haven't had to worry about simplifying too much. last year in precal, i would come in for extra help a lot, but i've always managed to only get B's in math :(

OpenStudy (amenah8):

so once i have my two x values, i plugged them into the equation to get my extremes (2.4 and 4.2) is that right?

OpenStudy (amenah8):

do i need a third point to determine the max and min?

OpenStudy (solomonzelman):

you have THREE critical number

OpenStudy (solomonzelman):

x=1 (when f'(X)=0) And when f'(x) -- undefined x=-1, 3

OpenStudy (solomonzelman):

Plug them into the function f(x), and tell me what do you get for each of them

OpenStudy (solomonzelman):

f(3) = ? f(1) = ? f(-1) = ?

OpenStudy (solomonzelman):

(give me exact answers please)

OpenStudy (amenah8):

when x = -1, y = 2.4 when x = 3, y = 4.2 when x = 1, y = 2.4

OpenStudy (solomonzelman):

that may not be correct.

OpenStudy (solomonzelman):

where x=-1 and x=3 solutions when we set the part in the root =0?

OpenStudy (solomonzelman):

So how come now you get different results for them now? \(\color{#000000 }{ \displaystyle f(x)=\sqrt{-x^2+2x+3} }\) \(\color{#000000 }{ \displaystyle f(-1)=\sqrt{-(-1)^2+2(-1)+3} =\sqrt{-1-2+3}=0 }\) \(\color{#000000 }{ \displaystyle f(3)=\sqrt{-3^2+2(3)+3} =\sqrt{-9+9}=0 }\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(1)=\sqrt{-1^2+2(1)+3} =\sqrt{4}=2 }\)

OpenStudy (amenah8):

isn't f(1) = swrt(6) ? --> -1^2=1, 2(1)=2, + 3 --> 1+2+3

OpenStudy (solomonzelman):

noo

OpenStudy (solomonzelman):

isn't f(1) = swrt(6) ? --> -1^2=1, 2(1)=2, + 3 --> -1+2+3

OpenStudy (amenah8):

:(

OpenStudy (solomonzelman):

-(1)² is still negative 1

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(1)=\sqrt{-1^2+2(1)+3} =\sqrt{-1+2+3}=\sqrt{4}=2 }\)

OpenStudy (solomonzelman):

there is a difference between (-1)² and -1²